OpenStudy (anonymous):

Which of the following is a solution to 2 radical (X-8=6) ?

OpenStudy (anonymous):

\[\sqrt{x-8}=6\] \[x-8=36\] \[x=44\]

OpenStudy (anonymous):

thats not one of the options :/

OpenStudy (anonymous):

what is the actual question? is it what i wrote or is it \[2\sqrt{x-8}=6\]?